# Towards Backdoor Stealthiness in Model Parameter Space

### v1

@@ 3. Comprehensive Backdoor Stealthiness
In this section, we first introduce the threat model (Section 3.1). Then, we analyze the behavior of neurons of backdoored models, suggesting that stealthy input- and feature-space backdoor attacks can be identified in parameter space (Section 3.2). Then, to achieve comprehensive stealthiness, we propose Grond that includes backdoor generation and Adversarial Backdoor Injection (Section 3.3).
@@ 3.1. Threat Model
Attacker’s capabilities. The attacker can train a well-performed surrogate model to generate UPGD, which is used to perturb the victim model’s input. Additionally, the attacker can alter the model’s weights during training. Table 12 in Appendix A.3 shows that the threat model of Grond is aligned with baseline attacks.
@@ 3.2. Lack of Parameter-Space Stealthiness
Taking this observation into consideration, we ran an initial experiment to understand the behavior of neurons of backdoored models. We use the TAC values (Zheng et al., 2022) to quantify the relevance of a neuron to the backdoor behavior according to the difference when the network accepts benign and backdoor inputs. A higher TAC value indicates that the neuron is strongly relevant for backdoor behaviors. Specifically, TAC is defined as:
@@ 3.3. Grond for Comprehensive Stealthiness
Similar to (Zeng et al., 2023; Zhang et al., 2021), UPGD is generated on a well-trained surrogate model trained on the clean training set. The architecture and parameters of the surrogate model do not necessarily need to be the same as the victim model (see Table 15 in Appendix B.2). UPGD is optimized following the PGD (Madry et al., 2018) algorithm to decrease the surrogate model’s cross-entropy loss that takes as inputs the adversarial examples (the poisoned samples in our case) and the target class label. This procedure is described formally in Algorithm 1. The $\bm{\delta}$ is the generated UPGD that will be used as a backdoor trigger; thus, $G_{x}({\bm{x}})={\bm{x}}+\bm{\delta}$. $S$ is the ball function with the radius $\epsilon$, and the small $\epsilon$ guarantees the imperceptibility of the backdoor trigger as it controls the perturbation’s magnitude.
@@
where $u$ is a fixed threshold and $\sigma$ is the UCLC value of the given weights. The measure for quantifying backdoor relevance can be changed from UCLC to others, such as the distance of neuron outputs when receiving benign and backdoor inputs, where a larger distance means the neuron is more relevant to backdoor behaviors and can be pruned. We use the modified UCLC for training efficiency as UCLC is data-free, which does not require calculation based on the outputs of neurons.
@@
Feature-space stealthiness. We hypothesize that feature-space stealthiness is a by-product of parameter-space and input-space stealthiness since the variation of feature maps is strongly correlated with model parameters and inputs. Figures 3 and 6 show that Grond can substantially increase the feature-space stealthiness. Detailed experimental analysis can be found in Section 4.2.
Table 2. Pruning-based mitigations against backdoored ResNet18 on CIFAR-10. BA refers to benign accuracy on clean data, ASR to attack success rate, and PR to the poisoning rate of the training set. The average drop of BA and ASR is also shown with downward arrows compared to the performance without any defense. Red marks indicate the attack failed to resist the defense with an ASR lower than 60%, and green means that the ASR is higher than 60%.
@@
BadNets (Gu et al., 2019) 93.13 100 92.42 71.71 91.6 1.06 88.99 49.02 84.04 13.82 89.26 $\downarrow$3.87 33.90 $\downarrow$66.1
@@
Narcissus (Zeng et al., 2023) 93.58 99.64 93.49 96.54 89.76 49.18 93.19 97.82 91.10 94.59 91.88 $\downarrow$1.7 84.53 $\downarrow$15.11
@@
Grond (PR=0.5%) 94.36 92.91 93.32 90.96 93.87 84.04 94.52 86.82 91.99 84.63 93.43 $\downarrow$0.93 86.61 $\downarrow$6.3
Table 3. Fine-tuning-based mitigations against backdoored ResNet18 on CIFAR-10.
@@
BadNets (Gu et al., 2019) 91.07 43.96 92.01 2.84 90.87 97.48 92.4 13.10 91.26 13.12 91.52 $\downarrow$2.06 34.10 $\downarrow$65.9
Blend (Chen et al., 2017) 91.64 99.61 92.52 1.73 91.84 8.84 93.4 100 91.86 100 92.25 $\downarrow$2.17 62.04 $\downarrow$37.96
WaNet (Nguyen and Tran, 2021) 91.11 0.99 90.89 1.03 87.98 0.81 92.17 0.04 90.3 4.89 90.49 $\downarrow$3.11 1.55 $\downarrow$97.82
IAD (Nguyen and Tran, 2020) 90.83 2.16 92.18 2.87 88.4 15.68 91.29 0.0 89.54 1.59 90.45 $\downarrow$2.43 4.46 $\downarrow$92.64
@@
Bpp (Wang et al., 2022b) 91.36 3.4 91.38 1.00 92.06 6.46 93.23 26.83 90.61 2.73 91.73 $\downarrow$2.46 8.08 $\downarrow$91.85
LC (Turner et al., 2019) 90.26 88.52 91.46 1.91 85.87 5.11 91.8 13.11 90.71 4.37 90.02 $\downarrow$4.29 22.60 $\downarrow$77.4
@@
Grond (PR=5%) 91.75 94.28 92.02 80.07 90.39 93.92 93.27 99.92 91.88 99.00 91.86 $\downarrow$1.57 93.44 $\downarrow$4.6
@@ 4. Experimental Evaluation
This section contains the main experimental results and analysis. Section 4.1 covers the datasets, baseline attacks, and defenses used in our experiments. Section 4.2 evaluates common backdoor attacks and Grond against pruning- and fine-tuning-based defenses. Section 4.3 provides an in-depth backdoor analysis, followed by Section 4.4, which explores how ABI can enhance common attacks. Section 4.5 covers backdoor model and input detection results, and Section 4.6 provides a comparison to supply-chain attacks. Finally, Section 4.7 covers the ablation study.
@@ 4.1. Experimental Setup
Datasets and Architectures. We follow the common settings in existing backdoor attacks and defenses and conduct experiments on CIFAR-10 (Krizhevsky et al., 2009), GTSRB (Stallkamp et al., 2012), and a subset of ImageNet (Deng et al., 2009) with 200 classes and 1,300 images per class (ImageNet200). More details about the datasets can be found in Appendix A.1. The primary evaluation is performed using ResNet18 (He et al., 2016). Moreover, we evaluate Grond using four additional architectures, VGG16 (Simonyan and Zisserman, 2015), DenseNet121 (Huang et al., 2017), EfficientNet-B0 (Tan and Le, 2019), and one recent architecture InceptionNeXt (Yu et al., 2024) (see Table 15 in Appendix B.2).
Attack Baselines. Grond is compared with 12 representative attacks: BadNets (Gu et al., 2019), Blend (Chen et al., 2017), WaNet (Nguyen and Tran, 2021), IAD (Nguyen and Tran, 2020), AdvDoor (Zhang et al., 2021), BppAttack (Wang et al., 2022b), LC (Turner et al., 2019), Narcissus (Zeng et al., 2023), Adap-Blend (Qi et al., 2023a), SSDT (Mo et al., 2024), DFST (Cheng et al., 2021), and DFBA (Cao et al., 2024). The default poisoning rate is set at 5% (of the training set) for all attacks following previous work (Xu et al., 2024a, b). Additionally, Grond is evaluated under various poisoning rates to provide a thorough analysis of its effectiveness. Following related works, the training schedule for attacks is 200 epochs when using CIFAR10 and GTSRB, and 100 epochs for ImageNet200. We use 1,000 images as the validation set to select the best-performing checkpoint. More implementation details and reasoning are provided in Appendix A.2.
Defense Baselines. We evaluate Grond and baseline attacks with 17 defenses, including four pruning-based methods (FP (Liu et al., 2018a), ANP (Wu and Wang, 2021), CLP (Zheng et al., 2022), and RNP (Li et al., 2023a)), five fine-tuning-based methods (vanilla FT, FT-SAM (Zhu et al., 2023), I-BAU (Zeng et al., 2022), FST (Min et al., 2023), and BTI-DBF(U) (Xu et al., 2024a)), five backdoor model detections (NC (Wang et al., 2019), Tabor (Guo et al., 2020), FeatureRE (Wang et al., 2022a), Unicorn (Wang et al., 2023), and BTI-DBF (Xu et al., 2024a)), two backdoor input detections (Scale-up (Guo et al., 2023) and IBD-PSC (Hou et al., 2024)), and a proactive defense CT (Qi et al., 2023b). Following their default settings, BTI-DBF (Xu et al., 2024a) and FP (Liu et al., 2018a) use 5% of training data, and other defenses use 1% of training data for detection or mitigation. CLP is a data-free backdoor pruning tool that uses no clean data. CT has access to the complete training set without knowing which samples are poisoned and is also able to interact with the model during training. Backdoor defense details and hyperparameters can be found in Appendix A.4.
@@ 4.2. Main Results on Backdoor Mitigation
All evaluated backdoor attacks are ineffective to at least one parameter-space backdoor defense on the CIFAR-10, as demonstrated in Tables 2 and 3. It suggests that common backdoor attacks designed to be stealthy in input and feature spaces are vulnerable to parameter-space defenses. Given that all backdoor behaviors are embedded in parameters of backdoored models, this finding suggests that future backdoor attacks should take parameter-space defenses into account as a standard step to evaluate comprehensive stealthiness.
Not surprisingly, Grond’s attack performance is better than all baseline attacks when considering evaluated backdoor defenses since Grond is designed to consider comprehensive stealthiness. On four pruning-based mitigations, Grond achieves 7.18% absolute higher ASR on average than the best backdoor attack, Narcissus. On five fine-tuning mitigations that show more powerful defense capability than pruning-based mitigations, Grond achieves 29.25% absolute higher ASR on average than Narcissus. In addition, Grond bypasses the five model detection and two input-space detections (see Section 4.5).
Pruning-based mitigation. We take a closer look at the details of pruning-based backdoor mitigation experiments in Table 2, presenting the results of all attacks against four pruning-based defenses. BadNets and Blend perform better on average than input-space stealthy attacks, e.g., WaNet and Bpp, because input-space stealthy attacks introduce significant separability in the feature space (see Figure 3). Across all pruning-based defenses, FP performs the worst, as expected, since it follows regular model pruning practice and is not a tailored backdoor pruning method.
@@
Figure 4. We show the performance of pruning neurons with high TAC values using different thresholds. The left column is BadNets, and the right is Grond.
@@ 4.3. Backdoor Analysis
Adaptive backdoor analysis by TAC pruning. Backdoor triggers are essential for calculating TAC values, making our TAC-based analysis highly adaptive to evaluating backdoor attacks. We directly utilize the trigger information to build a new pruning method based on the TAC analysis. In particular, we prune neurons with high TAC values in the backdoored model. Figure 4 shows the pruning results of Grond and BadNets, and TAC analysis for other attacks are in Appendix B.4. The first row of Figure 4 provides the pruning results. The second row contains the TAC values plots of neurons in the $4_{th}$ layer (the layer before the classification head) of ResNet18. For BadNets, the backdoor and benign behaviors of baseline attacks can be disentangled by pruning neurons with high TAC values. However, for Grond, pruning neurons with high TAC values decreases benign accuracy, which means the backdoor neurons are not easily distinguishable from benign neurons without harming benign performance. The analysis supports our statement that Grond spreads the backdoor to more neurons instead of a few prominent ones. We provide a sorted TAC value plot in Figure 9 in the appendix, showing the prominent neurons with high TAC values are quite limited in Grond.
Backdoor analysis by feature space inversion. We also provide a feature space analysis for different attacks by using a feature mask to decouple the benign and backdoor features following BTI-DBF (Xu et al., 2024a) and BAN (Xu et al., 2024b). The decoupling assumes that benign features related to the correct prediction introduce a lower loss, and the backdoor features related to backdoor prediction introduce a higher loss. In particular, the benign and inversed backdoor features are decoupled as follows:
where $g=f_{L-1}\circ\cdots f_{1}$, without the classification head. ${\bm{m}}$ is the learned feature mask, and $\mathcal{D}_{l}$ is a small set of benign samples with correct labels. As validated with BTI-DBF (Xu et al., 2024a) and BAN (Xu et al., 2024b), benign and backdoor features can be decoupled by the mask ${\bm{m}}$, after which backdoor features will introduce a substantially higher loss with respect to the ground truth label. More details can be found in Appendix A.6.
4.4. ABI Improves Common Backdoor Attacks
In this section, we show that our Adversarial Backdoor Injection (ABI) strategy generalizes to all evaluated common backdoor attacks. We combine the ABI module with baseline attacks to improve their resistance against parameter-space defenses. Figure 5 demonstrates that ABI is effective for all attacks when evaluating against the parameter-space defense ANP, where ASRs increase after adversarial injection, especially for BadNets, Blend, AdvDoor, Narcissus, and Adap-Blend. The improvement for feature space attacks (WaNet, IAD, and Bpp) is incremental. We speculate that feature space attacks rely too much on prominent features as their modification in the input space is minor. To activate the backdoor with such minor input modifications, the prominent features are required in the feature space. In addition, Figure 8 in Appendix B.3 shows the results of ABI without defense, demonstrating that it does not harm in general the BA and ASR when no defense is applied. Following our finding, we suggest that future backdoor attacks can use ABI to increase parameter-space stealthiness.
@@ 4.4. ABI Improves Common Backdoor Attacks
Table 5. Backdoor detection performance on CIFAR-10. 20 ResNet18 models are trained at each poisoning rate. Bd. refers to the number of models determined as backdoor models. Acc. refers to the detection accuracy.
@@
Table 6. Comparsion with supply-chain attacks.
Attack No Defense CLP (Zheng et al., 2022) FT-SAM (Zhu et al., 2023)
BA ASR BA ASR BA ASR
DFST (Cheng et al., 2021) 95.23 100 92.43 3.53 94.70 0.00
DFBA (Cao et al., 2024) 88.99 100 88.96 9.57 86.03 5.24
SSDT (Mo et al., 2024) 93.70 90.30 93.66 1.20 93.15 0.60
Grond 93.43 98.04 93.29 87.89 92.02 80.07

### v2

@@ 3.1. Threat Model
Attacker’s capabilities. The attacker can train a well-performed surrogate model to generate UPGD, which is used to perturb the victim model’s input. Additionally, the attacker can alter the model’s weights during training. Table 13 in App. B shows that the threat model of Grond is aligned with baseline attacks.
@@ 3.2. Lack of Parameter-Space Stealthiness
Considering this observation, we ran an initial experiment to understand the behavior of neurons in backdoored models. We use the TAC values (Zheng et al., 2022) to quantify the relevance of a neuron to the backdoor behavior. The TAC values show the change in the output of each neuron (a feature channel of convolutional layers) before and after the trigger is attached to the input. Thus, TAC quantifies a neuron’s sensitivity to the backdoor trigger. A high TAC value indicates that the neuron is strongly related to the backdoor behavior, whereas a low TAC value shows it is not. TAC takes the exact trigger information into account when measuring the backdoor effect, which makes TAC a straightforward and effective method that captures the relevant backdoor neurons’ behaviour. Specifically, TAC is defined as:
@@ 3.3. Grond for Comprehensive Stealthiness
Similar to (Zeng et al., 2023; Zhang et al., 2021), UPGD is generated on a well-trained surrogate model trained on the clean training set. The architecture and parameters of the surrogate model do not necessarily need to be the same as the victim model (see Table 15 in App. C.2). UPGD is optimized following the PGD (Madry et al., 2018) algorithm to decrease the surrogate model’s cross-entropy loss that takes as inputs the adversarial examples (the poisoned samples in our case) and the target class label. This procedure is described formally in Algorithm 1. The $\bm{\delta}$ is the generated UPGD that will be used as a backdoor trigger; thus, $G_{x}({\bm{x}})={\bm{x}}+\bm{\delta}$. $S$ is the ball function with the radius $\epsilon$, and the small $\epsilon$ guarantees the imperceptibility of the backdoor trigger as it controls the perturbation’s magnitude.
@@
where $u$ is a fixed threshold and $\sigma$ is the UCLC value of the given weights. The measure for quantifying backdoor relevance can be changed from UCLC to others, such as the distance of neuron outputs when receiving benign and backdoor inputs, where a larger distance means the neuron is more relevant to backdoor behaviors and can be pruned. We use the modified UCLC for training efficiency, as UCLC is data-free, which does not require calculation based on the outputs of neurons.
@@
Feature-space stealthiness. We hypothesize that feature-space stealthiness is a by-product of parameter-space and input-space stealthiness since the variation of feature maps is strongly correlated with model parameters and inputs. Figures 7 and 9 show that Grond can substantially increase the feature-space stealthiness.
Table 2. Pruning-based mitigations against backdoored ResNet18 on CIFAR10. BA refers to benign accuracy on clean data, ASR to attack success rate, and PR to the poisoning rate of the training set. The average drop of BA and ASR is also shown with downward arrows compared to the performance without any defense. Red marks indicate the attack failed to resist the defense with an ASR lower than 60%, and green means that the ASR is higher than 60%.
@@
BadNets (Gu et al., 2019) 93.13 100 92.42 71.71 91.60 1.06 88.99 49.02 84.04 13.82 89.26 $\downarrow$3.87 33.90 $\downarrow$66.10
@@
Narcissus (Zeng et al., 2023) 93.58 99.64 93.49 96.54 89.76 49.18 93.19 97.82 91.10 94.59 91.88 $\downarrow$1.70 84.53 $\downarrow$15.11
@@
DFST (Cheng et al., 2021) 95.23 100 94.79 93.84 94.64 3.72 92.43 3.53 93.29 12.68 93.79 $\downarrow$1.44 28.44 $\downarrow$71.56
DFBA (Cao et al., 2024) 88.99 100 86.85 0.03 88.96 9.55 88.96 9.57 88.96 0.90 88.43 $\downarrow$0.56 5.01 $\downarrow$94.99
@@
Grond (PR=0.5%) 94.36 92.91 93.32 90.96 93.87 84.04 94.52 86.82 91.99 84.63 93.43 $\downarrow$0.93 86.61 $\downarrow$6.30
Table 3. Fine-tuning-based mitigations against backdoored ResNet18 on CIFAR10.
@@
BadNets (Gu et al., 2019) 91.07 43.96 92.01 2.84 92.60 76.02 92.40 13.10 91.26 13.12 91.87 $\downarrow$1.26 29.81 $\downarrow$70.19
Blend (Chen et al., 2017) 91.64 99.61 92.52 1.73 91.84 8.84 93.40 100 91.86 100 92.25 $\downarrow$2.17 62.04 $\downarrow$37.96
WaNet (Nguyen and Tran, 2021) 91.11 0.99 90.89 1.03 87.98 0.81 92.17 0.04 90.30 4.89 90.49 $\downarrow$3.11 1.55 $\downarrow$97.82
IAD (Nguyen and Tran, 2020) 90.83 2.16 92.18 2.87 88.4 15.68 91.29 0.00 89.54 1.59 90.45 $\downarrow$2.43 4.46 $\downarrow$92.64
@@
Bpp (Wang et al., 2022b) 91.36 3.40 91.38 1.00 92.06 6.46 93.23 26.83 90.61 2.73 91.73 $\downarrow$2.46 8.08 $\downarrow$91.85
LC (Turner et al., 2019) 90.26 88.52 91.46 1.91 85.87 5.11 91.80 13.11 90.71 4.37 90.02 $\downarrow$4.29 22.60 $\downarrow$77.40
@@
SSDT (Mo et al., 2024) 93.74 0.70 93.15 0.60 90.27 3.10 92.85 0.20 90.79 1.40 92.16 $\downarrow$1.54 1.20 $\downarrow$89.10
DFST (Cheng et al., 2021) 95.01 2.07 94.70 0.00 89.75 19.11 93.06 2.66 90.41 22.34 92.59 $\downarrow$2.64 9.24 $\downarrow$90.76
DFBA (Cao et al., 2024) 86.68 10.10 86.03 5.24 85.48 100 82.76 57.62 84.49 100 85.09 $\downarrow$3.90 54.59 $\downarrow$45.41
Grond (PR=5%) 91.75 94.28 92.02 80.07 90.39 93.92 93.27 99.92 91.88 99.00 91.86 $\downarrow$1.57 93.44 $\downarrow$4.60
@@ 4.1. Experimental Setup
Datasets and Architectures. We follow the common settings in existing backdoor attacks and defenses and conduct experiments on CIFAR10 (Krizhevsky et al., 2009), GTSRB (Stallkamp et al., 2012), and a subset of ImageNet (Deng et al., 2009) with 200 classes and 1,300 images per class (ImageNet200). More details about the datasets can be found in App. A.1. The primary evaluation is performed using ResNet18 (He et al., 2016). Moreover, we evaluate Grond using four additional architectures, VGG16 (Simonyan and Zisserman, 2015), DenseNet121 (Huang et al., 2017), EfficientNet-B0 (Tan and Le, 2019), and one recent architecture InceptionNeXt (Yu et al., 2024) (see Table 15 in App. C.2). We also evaluate Grond with large and transformer-based models (see Table 16 in App. C.3).
Attack Baselines. Grond is compared with 12 representative attacks: BadNets (Gu et al., 2019), Blend (Chen et al., 2017), WaNet (Nguyen and Tran, 2021), IAD (Nguyen and Tran, 2020), AdvDoor (Zhang et al., 2021), BppAttack (Wang et al., 2022b), LC (Turner et al., 2019), Narcissus (Zeng et al., 2023), Adap-Blend (Qi et al., 2023a), SSDT (Mo et al., 2024), DFST (Cheng et al., 2021), and DFBA (Cao et al., 2024). The default poisoning rate is set at 5% (of the training set) for all attacks following previous work (Xu et al., 2024a, b). Additionally, Grond is evaluated under various poisoning rates to provide a thorough analysis of its effectiveness. Following related works, the training schedule for attacks is 200 epochs when using CIFAR10 and GTSRB, and 100 epochs for ImageNet200. We use 1,000 images as the validation set to select the best-performing checkpoint. More implementation details are provided in App. A.2.
Defense Baselines. We evaluate Grond and baseline attacks with 17 defenses, including four pruning-based methods (FP (Liu et al., 2018a), ANP (Wu and Wang, 2021), CLP (Zheng et al., 2022), and RNP (Li et al., 2023a)), five fine-tuning-based methods (vanilla FT, FT-SAM (Zhu et al., 2023), I-BAU (Zeng et al., 2022), FST (Min et al., 2023), and BTI-DBF (U) (Xu et al., 2024a)), five backdoor model detections (NC (Wang et al., 2019), Tabor (Guo et al., 2020), FeatureRE (Wang et al., 2022a), Unicorn (Wang et al., 2023), and BTI-DBF (Xu et al., 2024a)), two backdoor input detections (Scale-up (Guo et al., 2023) and IBD-PSC (Hou et al., 2024)), and a proactive defense CT (Qi et al., 2023b). Following their default settings, BTI-DBF (Xu et al., 2024a) and FP (Liu et al., 2018a) use 5% of training data, and other defenses use 1% of training data for detection or mitigation. CLP is a data-free backdoor pruning tool that uses no clean data. CT has access to the complete training set without knowing which samples are poisoned and can also interact with the model during training. Backdoor defense details and hyperparameters can be found in App. A.3.
@@ 4.2. Main Results on Backdoor Mitigation
All evaluated backdoor attacks are ineffective against at least one parameter-space backdoor defense on the CIFAR10, as demonstrated in Tables 2 and 3. It suggests that common backdoor attacks designed to be stealthy in input and feature spaces are vulnerable to parameter-space defenses. Given that all backdoor behaviors are embedded in parameters of backdoored models, this finding suggests that future backdoor attacks should consider parameter-space defenses as a standard step to evaluate comprehensive stealthiness.
Not surprisingly, Grond performs better than all baseline attacks when considering evaluated backdoor defenses since Grond is designed to consider comprehensive stealthiness. On four pruning-based mitigations, Grond achieves 7.18% higher ASR on average than the best backdoor attack, Narcissus. On five fine-tuning mitigations that show more powerful defense capability than pruning-based mitigations, Grond achieves 29.25% higher ASR on average than Narcissus. In addition, Grond bypasses the five model detection and two input-space detections (see Section 4.6).
Pruning-based mitigation. We take a closer look at the details of pruning-based backdoor mitigation experiments in Table 2, presenting the results of all attacks against four pruning-based defenses. BadNets and Blend perform better on average than input-space stealthy attacks, e.g., WaNet and Bpp, because input-space stealthy attacks introduce significant separability in the feature space (see Figures 7 and 9). Across all pruning-based defenses, FP performs the worst, as expected, since it follows regular model pruning practice and is not a tailored backdoor pruning method.
@@
Figure 3. Pruning neurons with high TAC values using different thresholds (the $x$ axis).
Figure 4. The left is the performance of pruning neurons with high TAC values using different thresholds for Grond. The right is the TAC analysis of Grond on 512 neurons.
Comparison with Supply-Chain Attacks. Sharing a similar threat model to supply-chain attacks, we compare Grond and three state-of-the-art supply-chain attacks (SSDT (Mo et al., 2024), DFST (Cheng et al., 2021), and DFBA (Cao et al., 2024)), where these attacks are also designed to be robust against backdoor defenses. In particular, DFST (Cheng et al., 2021) proposes to include a controlled detoxification technique in the training process, which restrains the model from picking up simple features. DFBA (Cao et al., 2024) directly modifies a few parameters of a classifier to inject a backdoor. SSDT (Mo et al., 2024) introduces additional terms in the loss for the Source-Specific and Dynamic-Triggers (i.e., SSDT) attack, which obscures the difference between normal samples and malicious samples. Tables 2 and 3 also include the performance of supply-chain attacks against pruning- and fine-tuning-based defenses. It is clear that existing backdoor defenses can defeat supply-chain attacks. The ASR of DFST (Cheng et al., 2021), DFBA (Cao et al., 2024), and SSDT (Mo et al., 2024) are decreased to less than 10% while the BA drop is less than 3%.
@@
4.3. Adaptive Defenses
As Grond includes UPGD trigger and Adversarial Backdoor Injection (ABI), we consider two adaptive defenses targeting those two components.
TAC pruning targeting the UPGD trigger. First, we use the UPGD trigger information to build a new pruning method based on the TAC values. The TAC values are calculated using the backdoor trigger, making the TAC highly adaptive to evaluating any backdoor attack. In particular, we prune neurons with high TAC values in the backdoored model, i.e., removing the neurons more sensitive to the UPGD trigger. Figure 4 shows the pruning results of Grond. The left figure provides the pruning results. The right figure contains the TAC values plots of neurons in the $4_{th}$ layer (the layer before the classification head) of ResNet18. We show that pruning neurons with high TAC values decreases benign accuracy, which means the backdoor neurons are not easily distinguishable from benign neurons without harming benign performance. The analysis supports our statement that Grond spreads the backdoor to more neurons instead of a few prominent ones. In App. C.7, we provide sorted TAC value plots in Figure 11, showing that prominent neurons with high TAC values are rather limited in Grond.
Neuron noise targeting ABI. Second, we consider adding noise to neurons of Grond models as an adaptive defense, as the ABI component involves operations on neuron weights. Specifically, we add noise to the weights and biases of batch normalization layers. The range of noise is limited by $[-\epsilon_{noise},\epsilon_{noise}]$. Table 5 provides the performance of Grond and the benign model under different noise levels. With increased noise, the benign accuracy of both benign and Grond models is decreased, but the ASR of Grond remains high. This result also supports the stealthiness of Grond in the parameter space.
@@ 4.4. Backdoor Analysis
This section analyzes why baseline attacks are ineffective against parameter space backdoor defenses, and why Grond performs better, according to the TAC values and the weights’ change after applying backdoor defenses.
Table 5. Adaptive defense against Grond using noise on neurons.
Method Benign Grond
BA ASR BA ASR
$\epsilon_{noise}=0.0$ 94.76 - 94.16 98.04
$\epsilon_{noise}=0.1$ 94.54 - 93.88 97.99
$\epsilon_{noise}=0.2$ 93.82 - 93.18 96.22
$\epsilon_{noise}=0.3$ 91.21 - 90.05 99.08
$\epsilon_{noise}=0.4$ 85.65 - 88.33 99.42
$\epsilon_{noise}=0.5$ 82.43 - 84.30 99.74
Pruning neurons with prominent TAC values for baseline attacks. Section 3.2 demonstrates the existence of prominent neurons with high TAC values in backdoor models. Note that TAC represents the strongest type of backdoor defense, where the exact trigger information is exploited. In this section, we show that pruning prominent neurons could mitigate all baseline attacks. Specifically, we assign zero to the neuron’s weight if its TAC value exceeds a certain threshold. Figure 3 shows the pruning performance using different thresholds for ten baseline attacks. We can observe that for all ten baseline attacks, one threshold can always be found where the ACC is high and the ASR is low. It indicates that the backdoor effect of the ten attacks can be erased while maintaining good performance on benign samples if these prominent neurons are pruned. Thus, we confirm the existence of prominent neurons, which corresponds to the backdoor effect. More importantly, these backdoor neurons can be disentangled from benign neurons, so these baseline attacks are not stealthy concerning neuron weights, i.e., not stealthy in the parameter space.
Grond’s weights are more difficult to be modified by backdoor defense. This section analyzes the changes while applying pruning- and fine-tuning-based defenses to baseline attacks and Grond. Our goal is to demonstrate that the baseline attacks can be significantly affected by defenses, but Grond can resist these defenses. In Figure 5, we record the changes in the weights of 512 neurons (in layer 4 of ResNet18) after applying two types of defenses, FT-SAM (Zhu et al., 2023) and CLP (Zheng et al., 2022), as they both focus on the backdoor-related neurons. In contrast to BadNets, the weight changes in Grond model after FT-SAM fine-tuning are smaller. In the sorted changes, it is clear that there are a few neurons for BadNets that correspond to significant changes, which is not the case for Grond. In the second row of Figure 5, the CLP can find a few neurons relevant to backdoor but cannot find these for Grond. In addition, pruning-based methods (as shown by CLP results in Figure 5) only improve a few neurons, while fine-tuning methods can update all neurons. We conjecture this is why fine-tuning-based defenses perform better than pruning-based defenses in Tables 2 and 3. More results with other baseline attacks are provided in Figure 12 in App. C.9.
Figure 5. The weight changes after backdoor defenses. More results with other attacks in Figure 12 in App. C.9.
@@
Table 6. Backdoor detection performance on CIFAR10. 20 ResNet18 models are trained at each poisoning rate. Bd. refers to the number of models determined as backdoor models. Acc. refers to the detection accuracy.
@@
4.5. ABI Improves Common Backdoor Attacks
In this section, we show that our Adversarial Backdoor Injection (ABI) strategy generalizes to all evaluated common backdoor attacks. We combine the ABI module with baseline attacks to improve their resistance against parameter-space defenses. Figure 6 demonstrates that ABI is effective for all attacks when evaluating against the parameter-space defense ANP, where ASRs increase after adversarial injection, especially for BadNets, Blend, AdvDoor, Narcissus, and Adap-Blend. The improvement for feature space attacks (WaNet, IAD, and Bpp) is incremental. We speculate that feature space attacks rely too much on prominent features, as their modification in the input space is minor. To activate the backdoor with such minor input modifications, the prominent features are required in the feature space. In addition, Figure 10 in App. C.6 shows the results of ABI without defense, demonstrating that it does not harm in general the BA and ASR when no defense is applied. Following our findings, we suggest that future backdoor attacks can use ABI to increase parameter-space stealthiness.